A 21.8-L tank of carbon dioxide (CO22) is at a pressure of 9.28 atm and temperature of 19.4∘∘C.

(a)Calculate the number of moles of gas in the tank.

______________ mol ( ± 0.02 mol)

(b)Obtain the number of grams of carbon dioxide in the tank. Hint: You can obtain the molar mass of CO22 by adding the molar mass of C and twice the molar mass of O. The molar mass of an element is given on the periodic table in units of grams/mole.

_____________ g ( ± 2 g)

(c)A fire breaks out, raising the ambient temperature to 69.1∘∘C while 82.3 g of gas leak out of the tank. Calculate the final pressure in the tank as a result of the fire and leakage. Hint: Use ?=R=0.0821 L⋅atmmol⋅KL⋅atmmol⋅K for the gas constant in the ideal gas law. This allows you to use liters for volume and atmospheres for pressure.

_____________atm ( ± 0.02 atm)
Answer #1

a)

V = Volume= 21.8 L = 0.0218 m3

P = pressure = 9.28 atm = 9.28 x 101325 Pa = 940296 Pa

n = number of moles

T = temperature = 19.4 C = 19.4 + 273 K = 292.4 K

Using the equation

PV = n RT

(940296)(0.0218) = n (8.314) (292.4)

n = 8.43 moles

b)

M = molar mass of CO2 = 44 g/mol

m = mass of CO2 = ?

mass is given as

m = n M

m = 8.43 x 44

m = 370.92 g

c)

m' = mass remaining = m - 82.3 g = 370.92 - 82.3 = 288.62 g

n' = number of moles = m'/M = 288.62/44 = 6.56

T = final temperature = 69.1 C = 69.1 + 273 K = 342.1 K

V = Volume= 21.8 L = 0.0218 m3

P = pressure

using the equation

PV = n' R T

P (21.8) = (6.56) (0.0821) (342.1)

P = 8.5 atm